Vinegar is approximately 4.5%(v/v) aqueous acetic acid [sp. gr. 1.05, purity=99.6% w/w]. How much NaOH would have to be added to 100mL of vinegar to completely neutralize the acid?
1
Expert's answer
2019-10-11T09:37:40-0400

V(vinegar)=100 mL

d(AcOH)=1.05 g/mL

x=0.045

"V(AcOH)=x*V(vinegar)"

"m(AcOH)=d(AcOH)*V(AcOH)=d(AcOH)*x*V(vinegar)"

"n(AcOH)=m(AcOH)\/M(AcOH)=d(AcOH)*x*V(vinegar)\/M(AcOH)"

"m(NaOH)=d(AcOH)*x*V(vinegar)*M(NaOH)\/M(AcOH)" "m(NaOH)=n(AcOH)*M(NaOH)" "n(NaOH)=m(NaOH)\/M(NaOH)=n(AcOH)"

m(NaOH)=d(AcOH)*x*V(vinegar)*M(NaOH)/M(AcOH)

m(NaOH)=1.05 (g/ml)*0.045*100 (ml)*40(g/mol)/ 60 (g/mol) = 3.15 g